# Verifying DNN-based Semantic Communication Against Generative Adversarial Noise

### v2

@@ 1 Introduction
Next-generation wireless networks are increasingly adopting semantic communication (SemCom) to address the growing demand for intelligent, task-oriented data transmission in applications such as autonomous vehicles, industrial IoT, and augmented reality [undefn, undefat, undefar]. While traditional communication systems focus on the technical level and treats wireless connectivity as a data pipe without regard for contextual meaning [undefau], SemCom prioritizes understanding the meaning behind transmitted messages and encodes only the necessary information to convey that meaning. In recent years, deep learning-based SemCom employs deep neural networks (DNNs) at both transmitter and receiver to transmit task-relevant semantic information, which addresses the evolving requirements of next-generation wireless applications [undefaj, undefn], e.g., adaptive compression under varying channel conditions, feature extraction for diverse task types, and real-time semantic understanding. Those are capabilities that traditional fixed coding schemes cannot provide but DNNs can learn through end-to-end optimization. In particular, DNN-based SemCom jointly trains source/channel coding and modulation alongside a model to perform a specific task. This helps eliminate the need to transmit and reconstruct full message.
Goal-oriented SemCom represents a paradigm shift from traditional bit-level transmission to meaning-focused communication, where the system optimizes for task-specific objectives (e.g., “pedestrian detected 10 meters ahead”) rather than accurately reconstructing the entire transmitted data, achieving up to 90% bandwidth reduction while preserving safety-critical decision making [undefat]. This approach is particularly crucial for emerging applications, such as autonomous vehicles [undefat] and industrial IoT [undefar], where the ultimate goal is to achieve specific tasks rather than perfect data recovery. DNNs excel in this domain [undefas, undefaf, undefag] because of their ability to learn complex semantic representations and optimize end-to-end performance across the entire communication pipeline, from source encoding/decoding to channel encoding/decoding and task execution.
@@
However, DNN-based SemCom is susceptible to adversarial attacks, in which small perturbations added to the inputs fooling DNNs to misclassify [undefa, undefy, undefaa, undefao, undefv]. Such attacks can have severe consequences in SemCom systems, e.g., in autonomous driving scenarios [undefat], adversarial perturbations could cause the system to misinterpret traffic signs (Fig. 1) or pedestrians, leading to potentially catastrophic safety failures. In industrial IoT applications [undefar], these attacks could result in incorrect sensor readings or control decisions, compromising the integrity of critical infrastructure. The consequences are particularly severe because semantic attacks target the meaning of transmitted data rather than just its reconstruction quality.
Existing attackers [undefab, undefa, undeff] focus on finding specific misclassification examples rather than establishing robustness guarantees across continuous input spaces. Recent probabilistic approaches [undefm, undefac] provide statistical guarantees for the safety of DNN models through hypothesis testing and safe region enumeration. However, these approaches still do not provide formal guarantees for the high-reliability DNN-based SemCom.
In this paper, we present the Formal Verification of Semantic Communication to Adversarial Noise (VScan) framework, which provides formal guarantees on the robustness of SemCom systems against bounded adversarial perturbations. Our key insight is that realistic adversarial constraints, e.g., power limitations and statistical undetectability requirements, can be encoded as logical formulae to enable efficient formal verification across multiple interconnected networks. The core technical challenge addressed by VScan is verifying properties across the complete SemCom pipeline from input to final output, e.g., through an encoder, an decoder, and a task-specific model, while handling multiple simultaneous noise sources including adversarial perturbations, input variations, and channel noise. This is challenging because traditional DNN verification focuses on single models with single perturbation sources, whereas SemCom requires coordinated analysis across three interconnected networks with multiple interacting noise sources that compound through the pipeline. By formulating the generative adversarial noise model, e.g., using MIP with sound over-approximated bounds, we enable the application of state-of-the-art DNN verification techniques [undefap, undefav, undefk] to verify robustness towards all adversarial noise and input perturbation within a continuous space in a sound and complete manner. More importantly, this work introduces SemCom as a new application domain for DNN verification, which is a relatively young field that has primarily focused on traditional adversarial robustness analysis [undefd], and brings mathematical rigor to wireless communication security with implications for safety-critical deployments.
@@ 2.1 DNN-based SemCom
DNN-based SemCom leverages deep learning’s universal function approximation capabilities to enable joint semantic-channel coding [undefc]. These systems employ end-to-end architectures with semantic encoders and decoders that extract and reconstruct meaning rather than exact bit sequences [undefas], introducing semantic-level metrics that better reflect communication effectiveness. For example, in autonomous vehicles, instead of transmitting high-resolution images, SemCom transmits concise messages detailing detected elements (e.g., “pedestrian 10 meters ahead”), significantly reducing bandwidth usage [undefat].
@@
Fig. 2 illustrates a practical DNN-based SemCom consisting of a transmitter (encoder $\mathcal{E}(\cdot)$), a wireless medium, and a receiver (decoder $\mathcal{D}(\cdot)$) [undefn]. The transmitter generates a compact semantic feature $z=\mathcal{E}(x)$ from the original data $x$, which is then transmitted over the wireless channel. Assuming perfect channel estimation, the receiver obtains a noisy signal $z^{\prime}=z+n$, where $n$ is additive white Gaussian noise (AWGN). The decoder recovers the message as $x^{\prime}=\mathcal{D}(z^{\prime})$, which is fed into a pragmatic model $\mathcal{F}(\cdot)$ to perform downstream tasks, yielding $y=\mathcal{F}(x^{\prime})$. The system is then jointly trained: the pragmatic model $\mathcal{F}(\cdot)$ first learns from original data $x$, then the encoder-decoder pair is optimized to extract and reconstruct essential semantic features that maximize $\mathcal{F}(\cdot)$ performance. This results in semantic features with a smaller footprint compared to traditional communication that aim for perfect reconstruction.
@@ 2.2 Adversarial Attacks
DNN models are vulnerable to a range of attacks targeting their accuracy [undefa, undefz]. Adversarial attack techniques aim to find an adversarial input, $\tilde{x}$, such that for a function, $F$, $F(\tilde{x})\neq F(x)$, where $x$ is the original input. Several methods exist for generating adversarial samples, e.g., Fast Gradient Sign Method (FGSM) [undefo] was designed to attack classification models that use Stochastic Gradient Descent [undef]. Given that FGSM relies on the model’s parameters, it is generally considered a white-box attack, indicating the attacker has full access to the model’s architecture and parameters. Note that a white-box attack can be extended to a black-box attack, in which the attacker builds surrogate models by sending queries to obtain input and output pairs from the system [undefy]. The Projected Gradient Descent (Pgd[undefab] is a more powerful multi-step variant of FGSM. Pgd operates on a schedule of iterative perturbations where the noise is clipped by a maximum allowed perturbation $\epsilon$ on each iteration.
@@ 2.3 Formal DNN Verification
Many DNN verifiers [undefs, undefap, undefk, undefh, undefi, undefj, undefg] treat the DNN verification problem as a satisfiability problem. More specifically, given a formula $\alpha$ representing an $L$-layer network $N$ with $N_{l}$ neurons in layer $l$:
@@
where actact is the activation function of the layer (e.g., ReLU, sigmoid, tanh), while $\phi_{in}\implies\phi_{out}$ represents the property to be proved. The DNN verification problem then can be reduced to a satisfiability problem [undefs, undefap, undefh, undefk, undefl]:
@@
The verifier returns unsat if Eq. 1 is unsatisfiable, indicating that $\phi$ is a valid property of $N$, and sat otherwise, indicating the $\phi$ is not a valid property of $N$.
State-of-the-art formal DNN verifiers from recent VNN-COMPs [undefd, undefan, undefb], such as $\alpha\beta$-Crown [undefav], NeuralSAT [undefk], Marabou [undefap], and PyRAT [undefw], employ different forms of branch-and-bound techniques to split verification problems into smaller subproblems and refine bounds through linear relaxations. These verifiers leverage CPU/GPU parallelization and advanced optimization techniques to efficiently handle large-scale networks. VScan leverages NeuralSAT and $\alpha\beta$-Crown, the two top tools at VNN-COMP’25 [undeft] as black-boxes and thus can easily integrate with other DNN verifiers.
@@ 4.1 Adversarial Attack on SemCom
An adversary injects malicious noise into the channel, directly corrupting transmitted semantic features and disrupting the DNN-based SemCom system (Fig. 2). We consider a white-box attack where the attacker knows the parameters of both the autoencoder and pragmatic model [undefa], generating adversarial noise to minimize expected system performance across various inputs (§4.2). White-box verification represents the worst-case scenario; if the system withstands this, it is robust against weaker black-box attacks [undefy]. We impose practical restrictions (§4.3, §4.4) ensuring the Pgm attack remains effective yet undetectable.
@@ 4.3 Power Constraint
Attackers must bound perturbation strength since overly strong noise is easily detected by energy monitors at the receiver. Unlike prior work using PSNR [undefy, undefx], we use peak noise ratio (PNR) which measures noise-to-signal power ratio in the semantic feature space (where attacks occur) rather than the image space. We impose the same magnitude limit $\rho$ per dimension during Pgm training:
@@ 4.4 Statistical Undetectability Constraint
Adversarial noise must be statistically indistinguishable from natural AWGN channel noise to evade pattern-based detection. The Pgm model [undefa] uses a GAN architecture [undefp] where discriminator $\mathcal{D}_{a}$ classifies whether latent space signals are Gaussian distributed. The generator $\mathcal{G}_{a}(\cdot)$ is trained to fool the discriminator, matching AWGN noise distribution:
@@ 5.2 Adversarial Perturbation Space
Computing the adversarial perturbation space generated by the generator model $\mathcal{G}_{a}(\cdot)$ of Pgm involves solving a mixed integer program (MIP) system [undefam]. This MIP formulation integrates signal power constraints that are specific to our SemCom verification task (§4.3).
@@
Note that the peak power constraint for each output dimension $\|y_{j}\|_{2}^{2}\leq\rho$ introduces quadratic constraints that standard abstraction-based verifiers cannot directly handle. To address this, we effectively employ an MIP solver, e.g., Gurobi [undefq], that supports quadratic programming to calculate these DNN output bounds under these power limitations. Note that, the selection of the MIP solver is not a limitation of VScan, as any MIP solver that supports quadratic programming can be integrated, e.g., CPLEX [undefr], SCIP [undefak].
@@ 5.4 Verification Framework
Adversarial AttackAlg. 1) attempts to find counterexamples on the composed network $\mathcal{F}\circ\mathcal{D}\circ\mathcal{E}$ and validates them against Pgm realizability to eliminate spurious results. Since counterexamples are drawn from the over-approximated bounds $n_{L},n_{U}]$ rather than the exact Pgm output space, they can be spurious. The validation step performs end-to-end attack from random trigger $r$ through the complete SemCom pipeline to confirm counterexample realizability. Finally, Formal Verification (Alg. 1) employs a DNN verifier on the complete system if attack methods fail. In particular, we employ two different state-of-the-art complete verifiers, NeuralSAT [undefk] and $\alpha\beta$-CROWN [undefav], to determine the satisfiability of generated verification problems.
@@ 6 Experimental Setups
Dataset.  We use two datasets widely adopted in SemCom literature [undefai, undefad, undefz]: (1) FashionMNIST [undefaq] with 28x28 grayscale images across 10 clothing categories; and (2) CIFAR10 [undefu] with 32x32 RGB images across 10 object classes. Both datasets compress features to 16-64 dimensions vs. original pixel space while preserving classification accuracy. The pragmatic model achieves 88.37% and 79.85% accuracy on FashionMNIST and CIFAR10, respectively.
@@
Tools.  We create two VScan variants: VScanABC uses $\alpha\beta$-CROWN [undefav] while VScanNSAT uses NeuralSAT [undefk]. VScan leverages these verifiers as black-box and can work with other DNN verifiers. We compare VScan against two adversarial attackers: (1) Pgm [undefa, undefz], an input-agnostic generative model that produces diverse perturbations without knowing transmitted content; and (2) Pgd [undefab, undefae], a gradient-based iterative attack using projected gradient descent with complete system knowledge. These cover input-agnostic generative attacks to input-specific optimization-based attacks.
@@ 8 Discussion
VScan provides formal analysis capabilities that are complementary to, rather than competitive with, existing empirical defense mechanisms. Adversarial training techniques [undefal] and defense methods like denoising autoencoders [undefah] focus on improving system robustness through training-time or inference-time modifications. In contrast, VScan performs post-hoc analysis of trained models to provide mathematical guarantees about their robustness properties. This distinction means that VScan can serve as a formal evaluation tool to rigorously assess whether adversarial training or other defense mechanisms have indeed improved model robustness.
@@ 9 Related Work
Recent years have witnessed a surge of interest in the security of deep learning-based semantic communication systems. Early work by Bahramali et al. [undefa] introduced robust adversarial attacks against DNN-based wireless communication, highlighting the vulnerability of such systems to perturbations generated by a well-trained generative network. Li et al. [undefy] extended this line of research by proposing black-box physical layer attacks, leveraging surrogate models and OFDM processing to enhance attack effectiveness. Other works have investigated the impact of channel statistics [undefao] and physical-layer characteristics [undefad] on both the efficacy of attacks and defenses. Hardware implementation on universal software radio peripheral and practical vulnerabilities have also been demonstrated [undefaa], underscoring the real-world relevance of these security concerns.
To defend against adversaraial attacks, DNN-based SemCom often incorporates denoising techniques, e.g., autoencoders [undefah], or model ensembling [undefaw]. However, these techniques significantly increase system complexity, e.g., running denoising steps to remove adversaraial noise before the SemCom. More importantly, they do not provide formal guarantees against the full spectrum of adversarial perturbations. Another popular mechanism is adversarial (re)training [undefal], which enhances the robustness of the target system. Yet, there is no formal guarantee that such defenses can prevent unseen attacks. More general, such defense mechanisms are developed based on a limited dataset of adversarial samples and leave systems vulnerable to perturbations outside the training set. Our work complements these approaches by providing a formal verification framework that mathematically proves SemCom’s robustness against adversarial attacks.

### v3

@@ 1 Introduction
Next-generation wireless networks are increasingly adopting semantic communication (SemCom) to address the growing demand for intelligent, task-oriented data transmission in applications such as autonomous vehicles, industrial IoT, and augmented reality [undefo, undefau, undefas]. While traditional communication systems focus on the technical level and treats wireless connectivity as a data pipe without regard for contextual meaning [undefav], SemCom prioritizes understanding the meaning behind transmitted messages and encodes only the necessary information to convey that meaning. In recent years, deep learning-based SemCom employs deep neural networks (DNNs) at both transmitter and receiver to transmit task-relevant semantic information, which addresses the evolving requirements of next-generation wireless applications [undefak, undefo], e.g., adaptive compression under varying channel conditions, feature extraction for diverse task types, and real-time semantic understanding. Those are capabilities that traditional fixed coding schemes cannot provide but DNNs can learn through end-to-end optimization. In particular, DNN-based SemCom jointly trains source/channel coding and modulation alongside a model to perform a specific task. This helps eliminate the need to transmit and reconstruct full message.
Goal-oriented SemCom represents a paradigm shift from traditional bit-level transmission to meaning-focused communication, where the system optimizes for task-specific objectives (e.g., “pedestrian detected 10 meters ahead”) rather than accurately reconstructing the entire transmitted data, achieving up to 90% bandwidth reduction while preserving safety-critical decision making [undefau]. This approach is particularly crucial for emerging applications, such as autonomous vehicles [undefau] and industrial IoT [undefas], where the ultimate goal is to achieve specific tasks rather than perfect data recovery. DNNs excel in this domain [undefat, undefag, undefah] because of their ability to learn complex semantic representations and optimize end-to-end performance across the entire communication pipeline, from source encoding/decoding to channel encoding/decoding and task execution.
@@
However, DNN-based SemCom is susceptible to adversarial attacks, in which small perturbations added to the inputs fooling DNNs to misclassify [undefa, undefz, undefab, undefap, undefw]. Such attacks can have severe consequences in SemCom systems, e.g., in autonomous driving scenarios [undefau], adversarial perturbations could cause the system to misinterpret traffic signs (Fig. 1) or pedestrians, leading to potentially catastrophic safety failures. In industrial IoT applications [undefas], these attacks could result in incorrect sensor readings or control decisions, compromising the integrity of critical infrastructure. The consequences are particularly severe because semantic attacks target the meaning of transmitted data rather than just its reconstruction quality.
Existing attackers [undefac, undefa, undeff] focus on finding specific misclassification examples rather than establishing robustness guarantees across continuous input spaces. Recent probabilistic approaches [undefn, undefad] provide statistical guarantees for the safety of DNN models through hypothesis testing and safe region enumeration. However, these approaches still do not provide formal guarantees for the high-reliability DNN-based SemCom.
In this paper, we present the Formal Verification of Semantic Communication to Adversarial Noise (VScan) framework, which provides formal guarantees on the robustness of SemCom systems against bounded adversarial perturbations. Our key insight is that realistic adversarial constraints, e.g., power limitations and statistical undetectability requirements, can be encoded as logical formulae to enable efficient formal verification across multiple interconnected networks. The core technical challenge addressed by VScan is verifying properties across the complete SemCom pipeline from input to final output, e.g., through an encoder, an decoder, and a task-specific model, while handling multiple simultaneous noise sources including adversarial perturbations, input variations, and channel noise. This is challenging because traditional DNN verification focuses on single models with single perturbation sources, whereas SemCom requires coordinated analysis across three interconnected networks with multiple interacting noise sources that compound through the pipeline. By formulating the generative adversarial noise model, e.g., using MIP with sound over-approximated bounds, we enable the application of state-of-the-art DNN verification techniques [undefaq, undefaw, undefl] to verify robustness towards all adversarial noise and input perturbation within a continuous space in a sound and complete manner. More importantly, this work introduces SemCom as a new application domain for DNN verification, which is a relatively young field that has primarily focused on traditional adversarial robustness analysis [undefd], and brings mathematical rigor to wireless communication security with implications for safety-critical deployments.
@@ 2.1 DNN-based SemCom
DNN-based SemCom leverages deep learning’s universal function approximation capabilities to enable joint semantic-channel coding [undefc]. These systems employ end-to-end architectures with semantic encoders and decoders that extract and reconstruct meaning rather than exact bit sequences [undefat], introducing semantic-level metrics that better reflect communication effectiveness. For example, in autonomous vehicles, instead of transmitting high-resolution images, SemCom transmits concise messages detailing detected elements (e.g., “pedestrian 10 meters ahead”), significantly reducing bandwidth usage [undefau].
@@
Fig. 2 illustrates a practical DNN-based SemCom consisting of a transmitter (encoder $\mathcal{E}(\cdot)$), a wireless medium, and a receiver (decoder $\mathcal{D}(\cdot)$) [undefo]. The transmitter generates a compact semantic feature $z=\mathcal{E}(x)$ from the original data $x$, which is then transmitted over the wireless channel. Assuming perfect channel estimation, the receiver obtains a noisy signal $z^{\prime}=z+n$, where $n$ is additive white Gaussian noise (AWGN). The decoder recovers the message as $x^{\prime}=\mathcal{D}(z^{\prime})$, which is fed into a pragmatic model $\mathcal{F}(\cdot)$ to perform downstream tasks, yielding $y=\mathcal{F}(x^{\prime})$. The system is then jointly trained: the pragmatic model $\mathcal{F}(\cdot)$ first learns from original data $x$, then the encoder-decoder pair is optimized to extract and reconstruct essential semantic features that maximize $\mathcal{F}(\cdot)$ performance. This results in semantic features with a smaller footprint compared to traditional communication that aim for perfect reconstruction.
@@ 2.2 Adversarial Attacks
DNN models are vulnerable to a range of attacks targeting their accuracy [undefa, undefaa]. Adversarial attack techniques aim to find an adversarial input, $\tilde{x}$, such that for a function, $F$, $F(\tilde{x})\neq F(x)$, where $x$ is the original input. Several methods exist for generating adversarial samples, e.g., Fast Gradient Sign Method (FGSM) [undefp] was designed to attack classification models that use Stochastic Gradient Descent [undef]. Given that FGSM relies on the model’s parameters, it is generally considered a white-box attack, indicating the attacker has full access to the model’s architecture and parameters. Note that a white-box attack can be extended to a black-box attack, in which the attacker builds surrogate models by sending queries to obtain input and output pairs from the system [undefz]. The Projected Gradient Descent (Pgd[undefac] is a more powerful multi-step variant of FGSM. Pgd operates on a schedule of iterative perturbations where the noise is clipped by a maximum allowed perturbation $\epsilon$ on each iteration.
@@ 2.3 Formal DNN Verification
Many DNN verifiers [undeft, undefaq, undefl, undefh, undefj, undefk, undefg] treat the DNN verification problem as a satisfiability problem. More specifically, given a formula $\alpha$ representing an $L$-layer network $N$ with $N_{l}$ neurons in layer $l$:
@@
where actact is the activation function of the layer (e.g., ReLU, sigmoid, tanh), while $\phi_{in}\implies\phi_{out}$ represents the property to be proved. The DNN verification problem then can be reduced to a satisfiability problem [undeft, undefaq, undefh, undefl, undefm]:
@@
The verifier returns unsat if Eq. 1 is unsatisfiable, indicating that $\phi$ is a valid property of $N$, and sat otherwise, indicating the $\phi$ is not a valid property of $N$ [undefi].
State-of-the-art formal DNN verifiers from recent VNN-COMPs [undefd, undefao, undefb], such as $\alpha\beta$-Crown [undefaw], NeuralSAT [undefl], Marabou [undefaq], and PyRAT [undefx], employ different forms of branch-and-bound techniques to split verification problems into smaller subproblems and refine bounds through linear relaxations. These verifiers leverage CPU/GPU parallelization and advanced optimization techniques to efficiently handle large-scale networks. VScan leverages NeuralSAT and $\alpha\beta$-Crown, the two top tools at VNN-COMP’25 [undefu] as black-boxes and thus can easily integrate with other DNN verifiers.
@@ 4.1 Adversarial Attack on SemCom
An adversary injects malicious noise into the channel, directly corrupting transmitted semantic features and disrupting the DNN-based SemCom system (Fig. 2). We consider a white-box attack where the attacker knows the parameters of both the autoencoder and pragmatic model [undefa], generating adversarial noise to minimize expected system performance across various inputs (§4.2). White-box verification represents the worst-case scenario; if the system withstands this, it is robust against weaker black-box attacks [undefz]. We impose practical restrictions (§4.3, §4.4) ensuring the Pgm attack remains effective yet undetectable.
@@ 4.3 Power Constraint
Attackers must bound perturbation strength since overly strong noise is easily detected by energy monitors at the receiver. Unlike prior work using PSNR [undefz, undefy], we use peak noise ratio (PNR) which measures noise-to-signal power ratio in the semantic feature space (where attacks occur) rather than the image space. We impose the same magnitude limit $\rho$ per dimension during Pgm training:
@@ 4.4 Statistical Undetectability Constraint
Adversarial noise must be statistically indistinguishable from natural AWGN channel noise to evade pattern-based detection. The Pgm model [undefa] uses a GAN architecture [undefq] where discriminator $\mathcal{D}_{a}$ classifies whether latent space signals are Gaussian distributed. The generator $\mathcal{G}_{a}(\cdot)$ is trained to fool the discriminator, matching AWGN noise distribution:
@@ 5.2 Adversarial Perturbation Space
Computing the adversarial perturbation space generated by the generator model $\mathcal{G}_{a}(\cdot)$ of Pgm involves solving a mixed integer program (MIP) system [undefan]. This MIP formulation integrates signal power constraints that are specific to our SemCom verification task (§4.3).
@@
Note that the peak power constraint for each output dimension $\|y_{j}\|_{2}^{2}\leq\rho$ introduces quadratic constraints that standard abstraction-based verifiers cannot directly handle. To address this, we effectively employ an MIP solver, e.g., Gurobi [undefr], that supports quadratic programming to calculate these DNN output bounds under these power limitations. Note that, the selection of the MIP solver is not a limitation of VScan, as any MIP solver that supports quadratic programming can be integrated, e.g., CPLEX [undefs], SCIP [undefal].
@@ 5.4 Verification Framework
Adversarial AttackAlg. 1) attempts to find counterexamples on the composed network $\mathcal{F}\circ\mathcal{D}\circ\mathcal{E}$ and validates them against Pgm realizability to eliminate spurious results. Since counterexamples are drawn from the over-approximated bounds $n_{L},n_{U}]$ rather than the exact Pgm output space, they can be spurious. The validation step performs end-to-end attack from random trigger $r$ through the complete SemCom pipeline to confirm counterexample realizability. Finally, Formal Verification (Alg. 1) employs a DNN verifier on the complete system if attack methods fail. In particular, we employ two different state-of-the-art complete verifiers, NeuralSAT [undefl] and $\alpha\beta$-CROWN [undefaw], to determine the satisfiability of generated verification problems.
@@ 6 Experimental Setups
Dataset.  We use two datasets widely adopted in SemCom literature [undefaj, undefae, undefaa]: (1) FashionMNIST [undefar] with 28x28 grayscale images across 10 clothing categories; and (2) CIFAR10 [undefv] with 32x32 RGB images across 10 object classes. Both datasets compress features to 16-64 dimensions vs. original pixel space while preserving classification accuracy. The pragmatic model achieves 88.37% and 79.85% accuracy on FashionMNIST and CIFAR10, respectively.
@@
Tools.  We create two VScan variants: VScanABC uses $\alpha\beta$-CROWN [undefaw] while VScanNSAT uses NeuralSAT [undefl]. VScan leverages these verifiers as black-box and can work with other DNN verifiers. We compare VScan against two adversarial attackers: (1) Pgm [undefa, undefaa], an input-agnostic generative model that produces diverse perturbations without knowing transmitted content; and (2) Pgd [undefac, undefaf], a gradient-based iterative attack using projected gradient descent with complete system knowledge. These cover input-agnostic generative attacks to input-specific optimization-based attacks.
@@ 8 Discussion
VScan provides formal analysis capabilities that are complementary to, rather than competitive with, existing empirical defense mechanisms. Adversarial training techniques [undefam] and defense methods like denoising autoencoders [undefai] focus on improving system robustness through training-time or inference-time modifications. In contrast, VScan performs post-hoc analysis of trained models to provide mathematical guarantees about their robustness properties. This distinction means that VScan can serve as a formal evaluation tool to rigorously assess whether adversarial training or other defense mechanisms have indeed improved model robustness.
@@ 9 Related Work
Recent years have witnessed a surge of interest in the security of deep learning-based semantic communication systems. Early work by Bahramali et al. [undefa] introduced robust adversarial attacks against DNN-based wireless communication, highlighting the vulnerability of such systems to perturbations generated by a well-trained generative network. Li et al. [undefz] extended this line of research by proposing black-box physical layer attacks, leveraging surrogate models and OFDM processing to enhance attack effectiveness. Other works have investigated the impact of channel statistics [undefap] and physical-layer characteristics [undefae] on both the efficacy of attacks and defenses. Hardware implementation on universal software radio peripheral and practical vulnerabilities have also been demonstrated [undefab], underscoring the real-world relevance of these security concerns.
To defend against adversaraial attacks, DNN-based SemCom often incorporates denoising techniques, e.g., autoencoders [undefai], or model ensembling [undefax]. However, these techniques significantly increase system complexity, e.g., running denoising steps to remove adversaraial noise before the SemCom. More importantly, they do not provide formal guarantees against the full spectrum of adversarial perturbations. Another popular mechanism is adversarial (re)training [undefam], which enhances the robustness of the target system. Yet, there is no formal guarantee that such defenses can prevent unseen attacks. More general, such defense mechanisms are developed based on a limited dataset of adversarial samples and leave systems vulnerable to perturbations outside the training set. Our work complements these approaches by providing a formal verification framework that mathematically proves SemCom’s robustness against adversarial attacks.